OpenStudy (anonymous):

Add and simplify. 3/4 + 1/10+7/15

OpenStudy (anonymous):

3/4 + 1/10 + 7/15 =45/60 + 6/10 + 28/60 =79 / 60

OpenStudy (anonymous):

so is that fraction in whole?

OpenStudy (radar):

It is an improper fraction. If you want it to be a mixed number, convert by dividing the numerator by the denominator getting 1 19/60.

OpenStudy (radar):

Do whatever the practice requires. They're both right.